OpenStudy (anonymous):

Solve for x: -2(x+3)=-2(x+1)-4

OpenStudy (anonymous):

you first need to simplify meaning get rid off the brackets

OpenStudy (anonymous):

\[-2x - 6 = 2x + 2 - 4\]

OpenStudy (anonymous):

did you try to solve this ? what answer did you get

OpenStudy (anonymous):

Im still doing it and two of the choices are a. All real numbers and b. no solution so what does that mean. The other choice are 2 and 3

OpenStudy (anonymous):

this is what i got: \[-2x - 6 = 2x + 2 - 4\] that is first to simplify! \[- 2 - 6 + 4 = 2x + 2x \]

OpenStudy (anonymous):

\[-4 = 4x\]

OpenStudy (anonymous):

\[x = -1\]

OpenStudy (anonymous):

so did you find a similar answer?